# Active Data Curation Effectively Distills Large-Scale Multimodal Models

### v1

@@ 4.4 Comparison to Prior Art
We now use our ACED method to pretrain models at large compute budgets, across three FLOP-scales, and compare with SoTA inference-efficient two-tower VLMs, including MobileCLIP [155], TinyCLIP [173], CLIP-KD [180], and CLIP-CID [183] (see supplementary for details). We further compare against the recently released DatologyAI CLIP models [3] (in grey), which have proprietary data curation strategies specific for classification and retrieval tasks independently. We train our ACED-F0, ACED-F1 and ACED-F2 models on the DataComp-1B dataset for 13B samples seen for a fair comparison. From Tab. 2, we observe that our ACED models are more FLOP-efficient and highly performant compared to other baselines in their respective classes—ACED-F0 outperforms MobileCLIP-S0 by $0.4$% and TinyCLIP-63M/32 by $1.9$%, on average, while being $10.81$% and $41.5$% more FLOP-efficient respectively; our ACED-F1 outperforms MobileCLIP-S1 by $1.8$%, and TinyCLIP-39M/16 by $10.2$%, on average, while being $6.5$% and $24.6$% more efficient respectively; ACED-F2 outperforms MobileCLIP-S2 by $1.1$% on average, while being $4.8$% more FLOP-efficient. Notably, on the most widely-used evaluations like ImageNet, COCO and Flickr, our method surpasses the previous SoTA by large margins—ACED-F0 outperforms MobileCLIP-S0 by $0.7$% on ImageNet, $1.5$% on COCO, and $2.8$% on Flickr; ACED-F1 outperforms MobileCLIP-S1 by $2.3$% on ImageNet, $2.6$% on COCO, and $4.7$% on Flickr, while ACED-F2 outperforms MobileCLIP-S2 by $2.5$% on ImageNet, $3.9$% on COCO, and $3.5$% on Flickr. In fact, our ACED-F1 model even outperforms MobileCLIP-S2 on ImageNet, while having $34$% lesser GFlops (see Fig. 1). This further validates the scalability of our ACED, especially given our models do not use any bespoke architectures or complex augmentations.
@@ 4.5 ACED yields better encoders for other tasks
We next evaluate the benefits of ACED specifically for training an auto-regressive text-decoder with a frozen image-encoder, in the LiT-Decoder setting [13]. We evaluate the trained LiT-decoder models on Flickr30k [116] captioning (using the CIDEr metric [156]) and visual-question-answering (using accuracy) tasks. Since prior works on these evaluation benchmarks use bigger foundation models (e.g., SigLIP [190] and SiLC [108]), we also train a larger ACED (B/16) model for 13B samples seen for fair comparison. Tab. 3 demonstrates that our ACED model outperforms both strong baselines across both tasks—particularly, our model outperforms competitors that have similar image-GFlops but are trained for a significantly higher number of samples seen (up to $\sim$3x). This further highlights the impact of ACED particularly for distilling knowledge in the image-encoders.
@@ 5 Conclusion
In this work we showed that active data curation implicitly implements a novel form of distillation, which combines knowledge from both a reference model and the data itself. With this insight, we developed ACID, a powerful method for distilling large multimodal encoders into much more efficient ones via online joint-example selection [35]. ACID strictly outperforms traditional forms of knowledge distillation in training contrastive VLMs. Given that ACID implicitly optimizes a different objective than traditional softmax-based KD, we further demonstrated these two objectives to be complementary, arriving at our final method, ACED, which combines the benefits of each. Using ACED we distilled models that set a new state-of-the-art for FLOP-efficient zero-shot classification and image-text reasoning.
@@ Appendix F About baselines and final ACED models
Appendix G Additional Experiments, Ablations and Results
@@
G.1 ACIDistill vs. IIDistill scaling
@@
G.2 Softmax vs Sigmoid Pretraining
@@
G.3 ACID vs KD as we scale compute
@@
G.4 Full Detailed Results across all 27 StableEval Evaluations
@@
G.5 Hyperparameter Sensitivity in ACID
@@
Appendix H Extended Related Works
Multimodal Data Curation. Recent works have emphasised the importance of data quality for pretraining multimodal models [109, 89, 48, 153, 41, 106]. Canonical methods for curating high-quality training data generally involve static offline curation include removing noisy samples [69, 20, 160, 144, 1, 19, 103, 178, 2, 48], rebalancing concept distributions [114, 178, 2], improving quality of text captions [91, 187, 82, 38, 110, 87, 88, 182, 192, 111], and using pretrained data-selector models for filtering samples with low image-text alignment [101, 169, 42, 168, 138, 139, 75, 186, 167].
@@ Appendix H Extended Related Works
Appendix I Discussion

### v2

@@ 4.4 Comparison to Prior Art
We now pretrain ACED models at large compute budgets, across three FLOP-scales, and compare with SoTA inference-efficient two-tower VLMs, including MobileCLIP [155], TinyCLIP [173], CLIP-KD [180], CLIP-CID [183] and proprietary DatologyAI-CLIP [3] (see Appendix F). We train our ACED-F0, ACED-F1 and ACED-F2 models on DataComp-1B for 13B samples seen. From Tab. 2, we observe that our ACED models are the most FLOP-efficient and highly performant—ACED-F0 outperforms MobileCLIP-S0 by $0.4$% and TinyCLIP-63M/32 by $1.9$%, on average, while being $10.81$% and $41.5$% more FLOP-efficient respectively; our ACED-F1 outperforms MobileCLIP-S1 by $1.8$%, and TinyCLIP-39M/16 by $10.2$%, on average, while being $6.5$% and $24.6$% more efficient respectively; ACED-F2 outperforms MobileCLIP-S2 by $1.1$% on average, while being $4.8$% more FLOP-efficient. Notably, on the most widely-used evaluations like ImageNet, COCO and Flickr, our method surpasses the previous SoTA by large margins—ACED-F0 outperforms MobileCLIP-S0 by $0.7$% on ImageNet, $1.5$% on COCO, and $2.8$% on Flickr; ACED-F1 outperforms MobileCLIP-S1 by $2.3$% on ImageNet, $2.6$% on COCO, and $4.7$% on Flickr, while ACED-F2 outperforms MobileCLIP-S2 by $2.5$% on ImageNet, $3.9$% on COCO, and $3.5$% on Flickr. In fact, our ACED-F1 model even outperforms MobileCLIP-S2 on ImageNet, while having $34$% lesser GFlops (see Fig. 1). This further validates the scalability of our ACED, especially given our models do not use any bespoke architectures or complex augmentations.
@@ 4.5 ACED yields better encoders for other tasks
We next evaluate the benefits of ACED specifically for training an auto-regressive text-decoder with a frozen image-encoder, in the LiT-Decoder setting [13]. We evaluate the trained models on Flickr30k [116] captioning (using CIDEr [156]) and visual-question-answering (using accuracy). Since prior works on these benchmarks use larger foundation models (e.g., SigLIP [190] and SiLC [108]), we also train a larger ACED (B/16) model for 13B samples seen Tab. 3 demonstrates that our ACED model outperforms both strong baselines across both tasks—particularly, our model outperforms competitors that have similar image-GFlops but are trained for a significantly higher number of samples seen (up to $\sim$3x). This further highlights the impact of ACED particularly for distilling knowledge in the image-encoders.
@@ 5 Conclusion
In this work we showed that active data curation implicitly implements a novel form of distillation, which combines knowledge from both a reference model and the data itself. With this insight, we developed ACID, a powerful method for distilling large multimodal encoders into much more efficient ones via online joint-example selection [35]. ACID strictly outperforms traditional forms of knowledge distillation in training contrastive VLMs. Given that ACID implicitly optimizes a different objective than traditional softmax-based KD, we further demonstrated these two objectives to be complementary, arriving at our final method, ACED, which combines the benefits of each. Using ACED we distilled models that set a new state-of-the-art for FLOP-efficient zero-shot classification and image-text retrieval.
@@ Appendix F About baselines and final ACED models
Appendix G Comparison with other batch selection methods
In this section, we compare our ACID method with other online batch selection methods in the literature as outlined in Sec. 2. For a fair comparison, we re-implement four batch-selection methods under our setting, namely, Bad-Students [34], Selective-Backprop [70], RHO-Loss [107] and JEST [35]. For this experiment, we pretrain SigLIP models on DataComp-1B [48] for 3B samples seen. For the reference models required by RHO-Loss, JEST and ACID, we use our pretrained WebLI-C++ reference.From Tab. 6, we observe that our ACID method outperforms all the other batch-selection methods by large margins ($1.4\%$ better than JEST and $3.2\%$ better than RHO-loss).
Method IN-val COCO 27-Avg
IID (baseline) 63.6 42.4 60.1
Softmax-KD 66.1 47.3 62.0
Bad-Students [34] 60.9 49.0 57.8
Sel-BP [70] 63.5 42.7 60.2
RHO-loss [107] 65.9 49.4 62.6
JEST [35] 68.7 53.4 64.4
ACID 71.0 53.6 65.8
Table 6: ACID outperforms all other online batch selection methods.
Appendix H Additional Experiments, Ablations and Results
@@
H.1 ACIDistill vs. IIDistill scaling
@@
H.2 Softmax vs Sigmoid Pretraining
@@
H.3 ACID vs KD as we scale compute
@@
H.4 Full Detailed Results across all 27 StableEval Evaluations
@@
H.5 Hyperparameter Sensitivity in ACID
@@
Appendix I Extended Related Works
Multimodal Data Curation. Recent works have emphasised the importance of data quality for pretraining multimodal models [109, 89, 48, 153, 41, 106]. Canonical methods for curating high-quality training data generally involve static offline curation include removing noisy samples [69, 20, 160, 144, 1, 19, 103, 178, 2, 48], rebalancing concept distributions [114, 178, 2], improving quality of text captions [91, 187, 82, 38, 110, 87, 88, 182, 192, 111], and using pretrained data-selector models for filtering samples with low image-text alignment [101, 169, 42, 168, 138, 139, 75, 186, 167]. Specifically, it has been shown that offline curation of noisy web-scale data can result in large pretraining efficiency gains [144, 69, 20, 160, 1, 19, 103, 178, 2, 101, 169, 42, 168, 138, 75, 186, 167, 159].
@@ Appendix I Extended Related Works
Appendix J Discussion on training cost vs baselines
In this section, we describe in detail the training costs required by ACID compared to other methods. We first define $F_{I}$ as the FLOPs-per-iteration of a forward pass of the image encoder of the student model. Similarly, we define $F_{T}$ as the FLOPs-per-iteration of a forward pass through the student text encoder. We do not consider the cost of the forward passes of teacher/reference models because we can cache their embeddings, as proposed in prior work [155, 155].
Given this, we compute the total FLOPs per normal IID iteration is 3($F_{I}$ + $F_{T}$). After caching reference embeddings, scoring the super-batch with the student model adds 4($F_{I}$ + $F_{T}$) for a filtering ratio of $0.8$, which gives a total FLOPs / iteration of 7($F_{I}$ + $F_{T}$) (7/3x overhead compared to IID training).
In Fig. 9, we show that the ACID method trained for 3B examples outperforms Softmax-KD training at 13B examples. Even with the 7/3x overhead, the absolute gains of using ACID are significant compared with additional IID and Softmax-KD training. Further, the main SoTA competition, MobileCLIP [155] has additional forward and backward passes due to an additional synthetic caption batch. This is an overhead of 3($F_{I}$ + $F_{T}$) - $F_{I}$ because the initial image forward-pass can be cached for the second batch. This gives a total FLOPs per iteration of 6($F_{I}$ + $F_{T}$) - $F_{I}$. If we compare for example, MobileCLIP-S0 (3.70 inference FLOPs) to ACED-F0 (3.30 inference FLOPs), the training per iteration of MobileCLIP-S0 = 6(2.39 + 1.32) - 2.39 = 19.81 FLOPs and ACED-F0 = 7(3.30) = 23.1 FLOPs. Thus ACED incurs an approx. 15% training overhead compared with MobileCLIP. However, it is worth noting that the methods proposed in Evans et al. [34] for flexible resolution scoring can be used to bring this training budget of ACED down drastically to well below that of MobileCLIP, with little loss in performance. We did not implement this as it has been shown in that prior work. Additionally, although MobileCLIP may have a slight training efficiency, their requirement for generating synthetic captions on new data is far more compute intensive than generating embeddings via our reference-model. Finally, we highlight that in general the main goal of our work (and others) is to maximize performance at given inference budgets as it is generally assumed that the training cost of efficient models will be amortized over model lifetime in use.
Appendix K Discussion